# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v1

@@ Abstract
Large Language Model (LLM)-based multi-agent systems (MAS) demonstrate remarkable potential for scientific discovery. Existing approaches, however, often automate scientific discovery using predefined workflows that lack rationality constraints. This often leads to aimless hypothesizing and a failure to consistently link hypotheses with evidence, thereby hindering systematic uncertainty reduction. Overcoming these limitations fundamentally requires systematic uncertainty reduction. We introduce PiFlow, an information-theoretical framework, treating automated scientific discovery as a structured uncertainty reduction problem guided by principles (e.g., scientific laws). In evaluations across three distinct scientific domains – discovering nanomaterial structures, bio-molecules, and superconductor candidates with targeted properties – our method significantly improves discovery efficiency, reflected by a 73.55% increase in the Area Under the Curve (AUC) of property values versus exploration steps, and enhances solution quality by 94.06% compared to a vanilla agent system. Overall, PiFlow serves as a Plug-and-Play method, establishing a novel paradigm shift in highly efficient automated scientific discovery, paving the way for more robust and accelerated AI-driven research. Code is publicly available at our GitHub.
@@ 1 Introduction
Large Language Model (LLM)-based Multi-Agent Systems (MAS) have significantly impacted automated scientific discovery Minaee2024LargeLM ; Wang2023ASO ; Zhang2024ACS across a wide range of fundamental fields, including chemistry Liu2024DrugAgentAA ; Ghafarollahi2024AtomAgentsAD ; Yang2024MOOSEChemLL ; Inoue2024DrugAgentED , biology Xiao2024CellAgentAL ; Nagarajan2025IANAI ; Averly2025LIDDIALI ; Ghafarollahi2024ProtAgentsPD , physics Jaiswal2024ImprovingPR , and material science Takahara2025AcceleratedIM ; Ghafarollahi2025AutomatingAD ; Ansari2024dZinerRI ; Wan2024FromTT ; Zhang2024LargeLM .
Although proficient in executing experiments within predefined workflows Lu2024TheAS ; lai2025prim , these systems often generate hypotheses that lack clear direction, leading to the uncertainty establishing clear links between hypotheses and their supporting or refuting evidence AI4Science2023TheIO ; Zhou2024HypothesisGW ; Baek2024ResearchAgentIR ; Schmidgall2025AgentLU ; Prabhakar2025OmniScienceAD ; Xie2023DARWINSD . Such a disconnect indicates inefficient exploration. Moreover, many of these approaches are tailored for specific tasks, often relying on meticulous prompt engineering that heavily incorporates domain knowledge (as detailed in Appendix A). Consequently, their ability to adapt to new scientific domains is often limited without substantial modifications Zhang2025ScientificLL ; Kumbhar2025HypothesisGF . These issues culminate in three primary challenges: (a) aimless hypothesizing; (b) unmaintained connections between hypotheses and evidence during exploration, hindering systematic validation and (c) limited generalization ability, where systems effective in one scenario (e.g., material science) often require substantial rework to be applicable in others.
To address these limitations, we introduce PiFlow, an information-theoretical framework for structured uncertainty reduction in scientific discovery. Viewing scientific exploration as a game against an unknown and challenging nature, where robust strategies are paramount, PiFlow employs Min-Max optimization: minimizing cumulative regret for exploitation, and maximizing information gain for hypothesis exploration. As a Plug-and-Play module, PiFlow integrates with MAS capable of hypothesizing and experimentation. Inspired by the challenge of navigating vast hypothesis spaces, PiFlow operates by using fundamental scientific principles, which may be initially proposed by or refined using LLMs, to steer the exploration.
The iterative selection of principles progressively reduces uncertainty in hypothesizing and in the interpretation of evidence, dynamically steering exploration by prioritizing those scientific principles that offer the highest instructive value for continued exploration. Figure 1 illustrates PiFlow’s method for assessing and utilizing scientific principles within a drug discovery context.
@@
This principle-aware approach yields systematic information gain: PiFlow selects high-potential principles and then guides hypothesizing via three actions, i.e., exploring, validating, or refining their scope and formulation. Thus, PiFlow progressively optimizes its guiding scientific principles to effectively steer hypothesizing. Furthermore, leveraging its Min-Max optimization, PiFlow theoretically achieves cumulative regret growth of $\mathcal{O}(\sqrt{T})$ over $T$ exploration steps. This sublinear regret underscores its guaranteed efficiency in navigating complex discovery landscapes. In summary, our contributions are:
@@
    We propose a novel paradigm for principle-aware scientific discovery, built upon an information-theoretical foundation that offers theoretical convergence guarantees.
@@
    We develop PiFlow, a Plug-and-Play framework that seamlessly integrates with existing MAS to enable focused exploration and exploitation, thereby enhancing discovery efficiency and operational flexibility.
@@
    We conduct extensive experiments across three distinct scientific discovery scenarios, demonstrating the broad applicability and significant performance improvements achieved by PiFlow.
@@ 2.1 Language Models for Scientific Discovery
Recently, large language models (LLM) have advanced scientific discovery with automation and rational design (Ren2025TowardsSI, ; Ma2024LLMAS, ). The internal knowledge of LLMs has demonstrated promising capability in focused chemical and material discovery (Yang2024MOOSEChemLL, ; Zhou2024HypothesisGW, ; pu2024leveraging, ; Ghafarollahi2024RapidAA, ). While tool-integrated LLMs like SciAgents (Ghafarollahi2024SciAgentsAS, ), DARWIN (Xie2023DARWINSD, ) and HoneyComb (Zhang2024HoneyCombAF, ) improve domain-specific reasoning and recall of factual insights, they still struggle to integrate physicochemical laws effectively when refining insights for design, risking biased proposals and inefficient exploration due to inherent hallucinations of LLMs (Zhang2023HowLM, ). Human-AI frameworks address this issue by leveraging the knowledge of domain experts (Reddy2024TowardsSD, ; Eythorsson2025TowardAS, ), yet remain limited to the scope of hypothesis generation (Alkan2025ASO, ), leading to insufficient exploration of complex chemical spaces (Luo2025LLM4SRAS, ). Surveys highlight persistent gaps in efficiency and interpretability (Zhang2024ACS, ; Han2024FromGT, ), underscoring the need for principled scientific discovery management beyond automated LLM reasoning (Ramos2024ARO, ).
@@ 2.2 Approaches of Multi-Agent Collaboration
Multi-agent systems (MAS) show promise for complex tasks (Lu2024TheAS, ; Ghafarollahi2024SciAgentsAS, ; Ni2023MechAgentsLL, ), yet their application to scientific discovery reveals limitations in current collaboration mechanisms (Tran2025MultiAgentCM, ). Rule-based methods (Zhang2023ExploringCM, ) offer consistency but their predefined rules lack the flexibility to incorporate nuanced scientific principles (e.g., physicochemical laws) or adapt to unexpected findings, hindering dynamic exploration. Role-playing approaches (Tran2025MultiAgentCM, ; He2024LLMBasedMS, ) leverage agent expertise, yet rigid roles can impede adaptation in scientific research (RamirezMedina2025AcceleratingSR, ; Lu2024TheAS, ; Ghafarollahi2024SciAgentsAS, ), and ensuring collective adherence to scientific principles when interpreting experimental insights is challenging. Model-based methods (Xu2023MAgICIO, ; Mu2023RuntimeVO, ; Li2023TheoryOM, ) aim for adaptability by learning from uncertainty, but struggle to build world models that accurately capture complex scientific phenomena and integrate guiding laws (Hao2023ReasoningWL, ), thereby impairing the balance between information perception and strategic, principle-guided reasoning.
Consequently, existing MAS paradigms often lack a dedicated awareness and systematic application of scientific principles during hypothesis generation and refinement (Gridach2025agentic, ; Luo2025LLM4SRAS, ; Reddy2024TowardsSD, ; Su2024ManyHA, ). This highlights a critical need for an explicitly principle-aware multi-agent collaboration framework. Our work addresses this gap, proposing a method where the collaborative discovery process is robustly guided by scientific principles to achieve more efficient and reliable outcomes.
@@ 3.1 Overview
We propose a principle-aware Multi-Agent System (MAS) designed to enhance scientific discovery via focused hypothesizing and structured exploration of hypothesis-evidence connections. Figure 2 illustrates the architecture. Its core comprises a Hypothesis-Validation loop that iteratively generates and tests hypotheses. PiFlow guides this loop by analyzing accumulated principle-outcome data. Strategic insights from PiFlow are relayed through a Planner agent ($\mathcal{A}_{P}$) to the Hypothesis Agent within the loop. Subsequent sections will elaborate on PiFlow’s specific architecture and its information-theoretical underpinnings.
@@ 3.2 Architecture
Our proposed principle-aware system leverages Large Language Model (LLM)-based MAS to conduct scientific discovery through strategic hypothesizing. The framework is comprised of two core, interconnected components: (1) an MAS that executes a Hypothesis-Validation loop, and (2) PiFlow, which serves as the strategic director for this loop.
@@
The scientific principles are foundational concepts, established laws, or patterns that explain phenomena within a specific scientific domain. These principles serve as high-level conceptual building blocks from which specific, testable hypotheses can be derived. This conceptualization aligns with broader discussions on the nature of scientific knowledge Poincar1906ScienceAH .
@@
    Hypothesis. Initially, a set of candidate principles $\mathcal{P}=\{p_{1},p_{2},p_{3}\}$ (see Definition 3.1)potentially proposed by domain experts or extracted by LLMs, is established. In each iteration $t$, $\mathcal{A}_{H}$ proposes a testable hypothesis, $h_{t}$, grounded in a selected principle $p_{i}\in\mathcal{P}$. The proposal is supported by structured rationales (comprising major and minor premises, following (Eisape2023ASC, )), rather than unconstrained association.
@@
    High-potential principle acquisition. After an initial phase of evidence collection, during which $\mathcal{T}_{t}$ is populated, PiFlow activates its core mechanism: an adversarial Min-Max optimization (detailed in Section 3.3). This optimization process analyzes $\mathcal{T}_{t}$ to identify a principle, $p^{*}$ (e.g., $p_{2}$ in Figure 2), predicted to have the highest potential for advancing the scientific inquiry.
@@
    Principle steering for hypothesizing. Following the identification of the highest-potential principle $p^{*}$ by the Min-Max optimization, PiFlow assigns a potential score to all principles in $\mathcal{P}$. This scoring enables a threshold-based partitioning detailed in Appendix C.1: principles with high scores (like $p^{*}$, e.g., $p_{2}$) drive refinement; those with medium scores trigger further validation ($p_{2}\rightarrow p_{2}^{\prime}$); and low scores prompt exploration of new conceptual areas.
This closed-loop feedback mechanism ensures that the Hypothesis-Validation cycle is continuously and adaptively steered by strategic insights derived from the system’s cumulative experience, as embodied in $\mathcal{T}_{t}$.
Plug-and-Play Modularity of PiFlow. The hypothesis steering mechanism above, which includes PiFlow and its interfacing Planner agent $\mathcal{A}_{P}$, has been intentionally engineered as a modular, Plug-and-Play system. This architectural choice ensures that principle-aware guidance can be readily integrated to enhance MAS engaged in scientific discovery as a part of prompts. The Planner agent, $\mathcal{A}_{P}$, serves as the critical adaptable interface, channeling analytically derived principle suggestions from PiFlow directly into the operational workflow of a host MAS. This adaptability positions PiFlow as a pivotal enhancement for automated scientific inquiry.
@@ 3.3 Min-Max Optimization in PiFlow
In our proposed method, strategic principle selection is achieved through a Min-Max optimization, as presented in Eq. 1. This approach is designed to balance the exploitation of established, high-potential principles (which then guide the formulation of specific hypotheses) with the exploration of novel ones, while explicitly incorporating information acquisition efficiency.
@@
Minimizing Cumulative Regret (Exploitation). The first term, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents the cumulative regret over $T$ iterations. Here, $v^{*}$ is a theoretical optimal outcome value, and $f^{*}(h_{t})$ is the outcome from hypothesis $h_{t}$. By minimizing this term, the policy $\pi$ is driven to exploit known high-potential principles and hypotheses to achieve outcomes as close as possible to the optimum $v^{*}$. This encourages the refinement and validation of promising avenues.
Maximizing Information Gain (Exploration). The second term $-\lambda\cdot I(h_{t};f^{*}|H_{t-1})$ promotes exploration. The policy $\pi$ seeks to minimize this term, which is equivalent to maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$. This mutual information quantifies the expected reduction in uncertainty about the true evaluation function $f^{*}$ upon observing the outcome of hypothesis $h_{t}$, given all past observations $H_{t-1}=\{\left(h_{m},y_{m}\right)\}_{m=1}^{t-1}$. The trade-off parameter $\lambda>0$ controls the balance between minimizing regret (exploitation) and maximizing information gain (exploration). A larger $\lambda$ places more emphasis on information acquisition.
@@
In summary, the Min-Max adversarial formulation underpinning PiFlow provides strong theoretical guarantees, notably sublinear regret bounds (formalized in Theorem 1, with proof in Appendix B). Importantly, PiFlow’s operational behavior, which involves practical approximations of this Min-Max solution, demonstrates consistent alignment with these theoretical expectations, as empirically validated in Section 5.3.

### v2

@@ Abstract
Large Language Model (LLM)-based multi-agent systems (MAS) demonstrate remarkable potential for scientific discovery. Existing approaches, however, often automate scientific discovery using predefined workflows that lack rationality constraints. This often leads to aimless hypothesizing and a failure to consistently link hypotheses with evidence, thereby hindering the systematic reduction of uncertainty. Overcoming these limitations fundamentally requires a principled approach to exploration. We introduce PiFlow, an information-theoretical framework, treating automated scientific discovery as a structured uncertainty reduction problem guided by principles (e.g., scientific laws). In evaluations across three distinct scientific domains – discovering nanomaterial structures, bio-molecules, and superconductor candidates with targeted properties – our method significantly improves discovery efficiency, reflected by a 73.55% increase in the Area Under the Curve (AUC) of property values versus exploration steps, and enhances solution quality by 94.06% compared to a vanilla agent system. Overall, PiFlow serves as a Plug-and-Play method, establishing a novel paradigm shift in highly efficient automated scientific discovery, paving the way for more robust and accelerated AI-driven research.
@@ 1 Introduction
Large Language Model (LLM)-based Multi-Agent Systems (MAS) have significantly impacted automated scientific discovery (Minaee et al., 2024; Wang et al., 2023; Zhang et al., 2024c) across a wide range of fundamental fields, including chemistry (Liu et al., 2024; Ghafarollahi & Buehler, 2024a; Yang et al., 2024c; Inoue et al., 2024), biology (Xiao et al., 2024; Nagarajan et al., 2025; Averly et al., 2025; Ghafarollahi & Buehler, 2024b), physics (Jaiswal et al., 2024), and material science (Takahara et al., 2025; Ghafarollahi & Buehler, 2025; Ansari et al., 2024; Wan et al., 2024; Zhang et al., 2024b).
Although proficient in executing experiments within predefined workflows (Lu et al., 2024; Lai & Pu, 2025), these systems often generate hypotheses that lack clear direction, leading to the uncertainty establishing clear links between hypotheses and their supporting or refuting evidence (AI4Science & Quantum, 2023; Zhou et al., 2024; Baek et al., 2024; Schmidgall et al., 2025; Prabhakar et al., 2025; Xie et al., 2023a). Such a disconnect indicates inefficient exploration. Moreover, many of these approaches are tailored for specific tasks, often relying on meticulous prompt engineering that heavily incorporates domain knowledge (as detailed in Appendix D). Consequently, their ability to adapt to new scientific domains is often limited without substantial modifications (Zhang et al., 2025; Kumbhar et al., 2025). These issues culminate in three primary challenges: (a) aimless hypothesizing; (b) unmaintained connections between hypotheses and evidence during exploration, hindering systematic validation and (c) limited generalization ability, where systems effective in one scenario (e.g., material science) often require substantial rework to be applicable in others.
To address these limitations, we introduce PiFlow, an information-theoretic framework for structured uncertainty reduction in scientific discovery. Viewing scientific exploration as a game against an unknown and challenging nature, where robust strategies are paramount, PiFlow employs Min-Max optimization: minimizing cumulative regret for exploitation, while maximizing information gain for efficient hypothesis exploration. As a Plug-and-Play module, PiFlow integrates with MAS capable of hypothesizing and experimentation. Inspired by the challenge of navigating vast hypothesis spaces, PiFlow operates by using fundamental scientific principles, which may be initially proposed by or refined using LLMs.
The iterative selection of principles progressively reduces uncertainty in hypothesizing and the interpretation of evidence, dynamically steering exploration by prioritizing those scientific principles that offer the highest instructive value for continued exploration. Figure 1 illustrates PiFlow’s method for assessing and utilizing scientific principles within a drug discovery context.
@@
This principle-aware approach yields systematic information gain: PiFlow selects high-potential principles and then guides hypothesizing via three actions, i.e., exploring, validating, or refining their scope and formulation. Thus, PiFlow progressively optimizes its guiding scientific principles to effectively steer hypothesizing. Furthermore, leveraging its Min-Max optimization, PiFlow theoretically achieves cumulative regret growth of $\mathcal{O}(\sqrt{T})$ over $T$ exploration steps (a detailed proof is provided in Appendix F). This sublinear regret underscores its guaranteed efficiency in navigating complex discovery landscapes. In summary, our contributions are:
@@
    We propose a novel paradigm for principle-aware scientific discovery, built upon an information-theoretical foundation that offers convergence guarantees.
@@
    We develop PiFlow, a Plug-and-Play framework that seamlessly integrates with existing MAS to enable focused exploration, thereby enhancing discovery efficiency and flexibility.
@@
    We conduct extensive experiments across three distinct scenarios, demonstrating the broad applicability and significant performance improvements achieved by PiFlow.
@@ 2.1 Language Models for Scientific Discovery
Recently, large language models (LLM) have advanced scientific discovery with automation and rational design (Ren et al., 2025; Ma et al., 2024). The internal knowledge of LLMs has demonstrated promising capability in focused chemical and material discovery (Yang et al., 2024c; Zhou et al., 2024; Pu et al., 2024; Ghafarollahi & Buehler, 2024c). While tool-integrated LLMs like SciAgents (Ghafarollahi & Buehler, 2024d), DARWIN (Xie et al., 2023a) and HoneyComb (Zhang et al., 2024a) improve domain-specific reasoning and recall of factual insights, they still struggle to integrate physicochemical laws effectively when refining insights for design, risking biased proposals and inefficient exploration due to inherent hallucinations of LLMs (Zhang et al., 2023b). Human-AI frameworks address this issue by leveraging the knowledge of domain experts (Reddy & Shojaee, 2024; Eythorsson & Clark, 2025), yet remain limited to the scope of hypothesis generation (Alkan et al., 2025), leading to insufficient exploration of complex chemical spaces (Luo et al., 2025). Surveys highlight persistent gaps in efficiency and interpretability (Zhang et al., 2024c; Han et al., 2024), underscoring the need for principled scientific discovery management beyond automated LLM reasoning (Ramos et al., 2024).
@@ 2.2 Approaches of Multi-Agent Collaboration
Multi-agent systems (MAS) show promise for complex tasks (Lu et al., 2024; Ghafarollahi & Buehler, 2024d; Ni & Buehler, 2023), yet their application to scientific discovery reveals limitations in current collaboration mechanisms (Tran et al., 2025). Rule-based methods (Zhang et al., 2023a) offer consistency but their predefined rules lack the flexibility to incorporate nuanced scientific principles (e.g., physicochemical laws) or adapt to unexpected findings, hindering dynamic exploration. Role-playing approaches (Tran et al., 2025; He et al., 2024) leverage agent expertise, yet rigid roles can impede adaptation in scientific research (Ramirez-Medina et al., 2025; Lu et al., 2024; Ghafarollahi & Buehler, 2024d), and ensuring collective adherence to scientific principles when interpreting experimental insights is challenging. Model-based methods (Xu et al., 2023; Mu et al., 2023; Li et al., 2023) aim for adaptability by learning from uncertainty, but struggle to build world models that accurately capture complex scientific phenomena and integrate guiding laws (Hao et al., 2023), thereby impairing the balance between information perception and strategic, principle-guided reasoning.
Consequently, existing MAS paradigms often lack a dedicated awareness and systematic application of scientific principles during hypothesis generation and refinement (Gridach et al., 2025; Luo et al., 2025; Reddy & Shojaee, 2024; Su et al., 2024). This highlights a critical need for an explicitly principle-aware multi-agent collaboration framework. Our work addresses this gap, proposing a method where the collaborative discovery process is robustly guided by scientific principles to achieve more efficient and reliable outcomes.
@@ 3.1 Overview
We propose a principle-aware MAS designed to enhance scientific discovery via focused hypothesizing and structured exploration of hypothesis-evidence connections. Figure 2 illustrates the architecture. Its core comprises a Hypothesis-Validation loop that iteratively generates and tests hypotheses. PiFlow guides this loop by optimizing accumulated principle-outcome data. Strategic insights dynamically optimized by PiFlow are relayed through a Planner agent ($\mathcal{A}_{P}$) to the Hypothesis Agent within the loop. Subsequent sections will elaborate on PiFlow’s specific architecture and its information-theoretical underpinnings.
@@ 3.2 Architecture
Our proposed principle-aware system leverages LLM-based MAS to conduct scientific discovery through strategic hypothesizing. The framework is comprised of two core, interconnected components: (1) an MAS that executes a Hypothesis-Validation loop, and (2) PiFlow, which serves as the strategic director for this loop.
@@
The scientific principles are foundational concepts, established laws, or patterns, articulated as natural language statements, that explain phenomena within a specific scientific domain. These principles serve as high-level conceptual building blocks from which specific, testable hypotheses can be derived. This conceptualization aligns with broader discussions on the nature of scientific knowledge (Poincaré, 1906).
@@
    Hypothesis. Initially, a set of dynamically growing candidate principles $\mathcal{P}=\{p_{1},p_{2},p_{3}\}$ (see Definition 3.1) potentially proposed by domain experts (see Appendix P) or extracted by LLMs, is established. In each iteration $t$, $\mathcal{A}_{H}$ proposes a testable hypothesis, $h_{t}$, grounded in a selected principle $p_{i}\in\mathcal{P}$. The proposal is supported by structured rationales (comprising major and minor premises, following Eisape et al. (2023), see Appendix T.2).
@@
    High-potential principle acquisition. After an initial phase of evidence collection, during which $\mathcal{T}_{t}$ is populated, PiFlow activates its core mechanism: an adversarial Min-Max optimization (detailed in Section 3.3). This optimization process analyzes $\mathcal{T}_{t}$ to identify a principle, $p^{*}$ (e.g., $p_{2}$ in Figure 2), predicted to balance the exploitation and exploration, i.e., have the highest potential, for advancing the scientific inquiry.
@@
    Principle steering for hypothesizing. Following the identification of the highest-potential principle $p^{*}$ by the Min-Max optimization, PiFlow assigns a potential score to all principles in $\mathcal{P}$. This scoring enables a threshold-based partitioning: principles with high scores (e.g., $p_{2}^{\prime}$) drive refinement; those with medium scores trigger further validation ($p_{2}\rightarrow p_{2}^{\prime}$); and low scores prompt exploration of new conceptual areas.
This closed-loop feedback mechanism ensures that the Hypothesis-Validation cycle is continuously and adaptively steered by strategic insights derived from the system’s cumulative experience, as embodied in $\mathcal{T}_{t}$. We provide a detailed analysis of the distinction from prompt engineering at Appendix E, and an illustrative example at Appendix H.
Plug-and-Play Modularity of PiFlow. The hypothesis steering mechanism above, which includes PiFlow and its interfacing Planner agent $\mathcal{A}_{P}$, has been intentionally engineered as a modular, Plug-and-Play system. This architectural choice ensures that principle-aware guidance can be readily integrated to enhance MAS engaged in scientific discovery as a part of prompts. As demonstrated in Appendix K, we successfully integrate PiFlow with existing ChemToolAgent (Yu et al., 2024) without any architecture modifications. This adaptability positions PiFlow as a pivotal enhancement for automated scientific inquiry.
@@ 3.3 Min-Max Optimization in PiFlow
In our proposed method, strategic principle selection is achieved through a Min-Max optimization, as presented in Eq. 1. This approach is designed to balance the exploitation of established, high-potential principles (which then guide the formulation of specific hypotheses) with the exploration of novel ones, while explicitly incorporating information acquisition efficiency:
@@
Minimizing cumulative fegret (exploitation). The first term, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents the cumulative regret over $T$ iterations. Here, $v^{*}$ is a theoretical optimal outcome value, and $f^{*}(h_{t})$ is the outcome from hypothesis $h_{t}$. By minimizing this term, the policy $\pi$ is driven to exploit known high-potential principles and hypotheses to achieve outcomes as close as possible to the optimum $v^{*}$. This encourages the refinement and validation of promising avenues.
Maximizing information gain (exploration). The second term $\lambda\cdot I(h_{t};f^{*}|H_{t-1})$ promotes exploration. The policy $\pi$ seeks to minimize this term, which is equivalent to maximizing the mutual information $I(h_{t};f^{*}|H_{t-1})$. This mutual information quantifies the expected reduction in uncertainty about the true evaluation function $f^{*}$ upon observing the outcome of hypothesis $h_{t}$, given all past observations $H_{t-1}=\{\left(h_{m},y_{m}\right)\}_{m=1}^{t-1}$. The trade-off parameter $\lambda>0$ controls the balance between minimizing regret (exploitation) and maximizing information gain (exploration). A larger $\lambda$ places more emphasis on information acquisition.
@@
Building upon the theoretical framework above, we derive a computationally tractable algorithm (Algorithm 1) that serves as a principled approximation of the abstract Min-Max objective in Eq. 1. The full rationale and derivation for this approximation are detailed in Appendix G.
In summary, the Min-Max adversarial formulation underpinning PiFlow provides strong theoretical guarantees, notably sublinear regret bounds (formalized in Theorem 1, with proof in Appendix F). Importantly, its operational behavior, which involves practical approximations of this Min-Max solution, demonstrates consistent alignment with these theoretical expectations, as empirically validated in Section 5.4.